# Atomic Smart Contract Interoperability with High Efficiency via Cross-Chain Integrated Execution

### v1

@@ Abstract
Index Terms:
Blockchain Interoperability, Cross-Chain Integrated Execution, Efficient and Atomic Interoperability Protocol, Cross-Chain Smart Contract Invocation.
@@ I Introduction
With the advent of Bitcoin and Ethereum [1, 2], we have witnessed the emergence of an increasing number of programmable blockchains  [3, 4, 5, 6]. On these programmable blockchains, developers can write and deploy smart contracts to build various complex decentralized applications (dApps, such as DeFi, NFT, etc. [7, 8]). Among these programmable blockchains, those that share compatible smart contract execution environment with Ethereum (i.e., Ethereum Virtual Machine, a.k.a. EVM [9]) dominate the landscape, accounting for over 90% of the total value locked [10]. With the increasing number of EVM-compatible blockchains and the diversity of dApps running on each chain, the demand for cross-chain dApps has grown significantly [11]. Cross-chain dApps refer to dApps that require cross-chain smart contract invocations (CCSCI) and coordinated execution across multiple blockchains [12].
However, ensuring overall atomicity for the entire cross-chain dApp while efficiently handling CCSCI remains a critical challenge. As illustrated in Figure 1 left, consider a classic train-and-hotel problem [13]. A user wants to book an outbound train ticket (on Train Chain) through a travel agency (on Agency Chain), then book a hotel (on Hotel Chain), followed by a return train ticket (again on Train Chain). The user wants to ensure that the entire series of CCSCI either all succeed or all fail, ensuring overall atomicity. More importantly, maintaining efficiency during CCSCI is crucial. Key considerations include how to minimize latency, reduce gas consumption (monetary cost), and improve concurrency during the CCSCI process. More cross-chain dApp scenarios are discussed in Section VIII.
@@
The design of IntegrateX faces several challenges, which we will address below.
Challenge 1: How to efficiently and securely deploy smart contracts across chains to the same blockchain. We address this by proposing an on-chain/off-chain Hybrid Cross-Chain Smart Contract Deployment Protocol. Specifically, to reduce the overhead associated with cross-chain deployment of smart contracts (thereby improving efficiency), we devise a set of guidelines that allow developers to decouple smart contracts into logic execution contracts and state storage contracts. This approach enables our protocol to clone and deploy only the logic contracts onto the same target blockchain, while the state-heavy contracts remain in their original locations. To further reduce the gas consumption during cross-chain deployment, we propose to clone and deploy the logic contracts to the same blockchain via an off-chain mechanism. However, malicious actors may tamper with the contracts during the off-chain clone and deployment process, leading to compromised security. To address this, we design and deploy bridging smart contracts on each blockchain. These bridging smart contracts perform on-chain cross-chain comparison and verification of the logic contracts between the source chain and the target chain to ensure security.
Challenge 2: How to enhance concurrency and reduce overhead during CCSCI while ensuring overall atomicity. We address this by proposing a Cross-Chain Smart Contract Integrated Execution Protocol. Specifically, to ensure overall atomicity during CCSCI, we adopt a 2PC-based mechanism, similar to existing mainstream approaches. This process involves locking the relevant states across the involved chains, transmitting the states to the chain responsible for integrated execution, executing all logic, and then returning the states to the respective chains to unlock and update them. This guarantees that the cross-chain dApp either completes entirely or is fully rolled back. Moreover, cross-chain state transfers may incur significant overhead. To mitigate the overhead of cross-chain communication, we aggregate multiple cross-chain transactions into a single one where necessary, and transmit them across chains. More importantly, a straightforward state-locking mechanism reduces transaction concurrency. To enhance concurrency, we establish a set of guidelines that allow developers to further decompose and lock contract states at a finer granularity. By locking more granular states, our protocol alleviates the issue of poor concurrency that arises when entire states are locked.
@@ II Background and Related Work
There are few other solutions that guarantee overall atomicity for the cross-chain dApp with complex CCSCI. However, they generally suffer from inefficiency  [23, 24, 25]. A complex cross-chain dApp typically involves multiple blockchains and several interdependent application logic components (implemented through smart contracts). In these works, handling such complexity requires multiple rounds of cross-chain execution and cross-chain interaction in sequential order across several involved chains. Specifically, this involves locking relevant states, executing the logic fragment, reaching consensus, transferring intermediate states to the next blockchain responsible for executing the subsequent logic fragment, and finally unlocking and updating the states. As a result, they typically suffer from high latency and poor concurrency handling CCSCI. Additionally, Atomic IBC [24] relies on the Cosmos Hub to provide security guarantees for its protocol. This approach is not easily adaptable to other ecosystems (e.g., EVM-compatible blockchains) and may introduce security bottlenecks.
@@ III System Overview
In this section, we provide an overview of IntegrateX. IntegrateX’s architecture and workflow overview is shown in Figure 2. For a detailed analysis of IntegrateX’s security and further discussion, please refer to Section VI and Section VIII.
III-A System and Threat Model
In IntegrateX, there are $n$ (variable) blockchains that share the same smart contract execution environment (e.g., EVM). These blockchains are developed by their respective projects (e.g., Ethereum, BSC, etc. [2, 46]) and operated by their own blockchain nodes, which handle transaction processing, consensus within the blockchain, and other tasks. IntegrateX also features $m$ (variable) relayers responsible for trustless cross-chain communication between blockchains, similar to many mainstream interoperability protocols [24, 17, 47, 48]. Each relayer has its own public-private key pair and signs cross-chain transactions. Additionally, IntegrateX’s bridging smart contracts are deployed on each blockchain (similar to on-chain light clients [17, 24]). The bridging contracts mainly serve to verify state transitions on the other blockchains, and send information externally via event emissions. The above IntegrateX’s transport layer architecture is similar to some well-known cross-chain communication protocols (e.g., IBC) [17, 24], which guarantee basic security during cross-chain communication. However, IntegrateX achieves efficiency and overall atomicity at the application layer, which these other protocols do not.
IntegrateX also includes several intra-chain dApp providers. These providers can freely choose to deploy their contracts and run their intra-chain dApps on different blockchains. Additionally, there are cross-chain dApp providers who can similarly flexibly choose to deploy their cross-chain dApps across various blockchains. A cross-chain dApp typically consists of multiple intra-chain dApps distributed across different blockchains, and relies on IntegrateX’s cross-chain protocol to ensure efficient CCSCI and maintain the overall atomicity of the cross-chain dApp. Finally, there are also users within the IntegrateX ecosystem. These users interact with the cross-chain dApps by sending transactions to the blockchains in order to use these cross-chain dApps.
@@ III-A System and Threat Model
Threat Model. In IntegrateX, we make only minimal trust assumptions about the relayers, assuming that at least one relayer is honest and functioning correctly. This assumption is consistent with those made in many existing secure interoperability protocols [24, 17, 47, 48]. For each blockchain, the proportion of Byzantine nodes is assumed to be less than the fault tolerance threshold $t$ of the respective blockchain network (e.g., for blockchains using BFT-type consensus protocols under partial-synchronous network, $t=1/3$ [49]). This ensures the safety and liveness of consensus within each blockchain. As for the dApp providers and users—who represent the application layer components—we make no specific threat assumptions, similar to most existing works [17, 25, 23]. However, we do discuss common countermeasures for dealing with malicious behavior from these components in Section VIII.
III-B Objective
@@
Figure 2: An overview of IntegrateX.
III-C Primary Workflow
@@
The operation of IntegrateX consists of three main phases: Smart Contract Preparation, Cross-Chain Smart Contract (CCSC) Deployment, and Cross-Chain Smart Contract (CCSC) Integrated Execution, as illustrated in Figure 2.
Smart Contract Preparation. This phase only occurs when dApp providers need to develop or upgrade smart contracts, so the frequency of this phase is relatively low. In this phase, dApp providers can flexibly develop logic contracts and state contracts according to our defined logic-state decoupling guidelines (Section IV-A). This decoupling is primarily intended to reduce gas consumption and minimize storage costs on the target chain by cloning only the logic contracts during the subsequent CCSC deployment phase. Additionally, dApp providers can develop their contracts according to our fine-grained state lock guidelines (Section V-C), which enhances transaction concurrency during the CCSC integrated execution phase. Once the smart contract development is complete, dApp providers can freely choose a blockchain to deploy their contracts.
CCSC Deployment. This phase only runs when cross-chain dApp providers need to deploy logic contracts across chains onto the same blockchain, which also occurs infrequently. In this phase, cross-chain dApp providers can flexibly choose a target chain for deployment. By sending a transaction, they can clone and deploy the specified logic contracts from other designated chains onto the chosen target chain based on their cross-chain dApp needs. IntegrateX introduces an off-chain clone approach (Section IV-B) to reduce gas consumption, while on-chain cross-chain verification (Section IV-C) ensures that the cloned contracts are identical to the original ones.
CCSC Integrated Execution. This phase runs every time a user needs to interact with a cross-chain dApp, making it the most frequent phase of operation. In this phase, users interact with cross-chain dApps by sending transactions to the target chain, invoking cross-chain smart contracts based on the dApp’s application logic. IntegrateX employs an atomic integrated execution scheme (Section V-A), similar to the 2PC protocol, to ensure the overall atomicity of the series of CCSCI involved in a cross-chain dApp. This scheme first locks the relevant states of the smart contracts involved in the cross-chain dApp on the respective chains at a fine-grained level, then transmits these states across chains to the target chain. Since the target chain already contains all the execution logic required for the cross-chain dApp (from the earlier phase), it can integrate and execute all logic within one transaction once the necessary states have been received. After execution, the new states are returned to the corresponding chains for unlocking and updating. Additionally, IntegrateX employs a transaction aggregation mechanism (Section V-B) during this phase to reduce cross-chain overhead.
In the subsequent descriptions within this paper, we will refer to the target chain selected by a cross-chain dApp for deployment—i.e., the chain responsible for integrated execution—as the execution chain. The other chains associated with the cross-chain dApp will be collectively referred to as invoked chains.
@@ IV-A1 Logic-State Decoupling Example
function UpdateSteate(bytes[] memory args) public returns();
@@
function UpdateSteate(bytes[] memory args) public returns();
@@
(a) Off-chain clone and deployment
(b) On-chain verification
Figure 3: Hybrid Cross-Chain Smart Contract Deployment Protocol.
@@ IV-B Off-Chain Clone and Deployment
Transmitting contract bytecode on-chain (via contract event) requires a significant amount of gas. To improve the efficiency of cross-chain logic deployment and reduce gas consumption, we propose transferring the contract bytecode via an off-chain solution. The off-chain clone and deployment consists of two phases: the preparation phase and the clone phase. The main process is shown in Figure  3(a).
Preparation. In this phase, the cross-chain dApp provider needs to obtain the call tree of smart contracts and determine which logic contracts need to be cloned. This can be achieved by using some static analysis tools [50] and other methods, similar to existing works [51, 23, 25]. After this, the developer can choose a blockchain to deploy the cross-chain dApp. It is important to note that we offer developers a high degree of flexibility: They can select any blockchain according to their preference. To reduce costs, they may also choose a blockchain that already contains some of the invoked logic contracts. Once the selection is made, the developer sends a transaction to invoke the bridging contract on the chosen chain. The bridging contract then triggers an event to notify the relayers to initiate the cross-chain deployment. The event includes the ID of the invoked chain and the addresses of the logic contracts $Addr_{\text{L}}$ to be cloned on the invoked chain. This concludes the preparation phase.
@@
Remarks. For the cross-chain reliability of the protocol, multiple relayers perform the clone process after detecting the bridging contract’s event. Therefore, even if some relayers do not respond to the event, other relayers will ultimately complete the contract clone and deployment. Once one relayer completes the deployment, the bridging contract will trigger an event to stop the others, ensuring that the logic contract will not be deployed multiple times.

### v2

@@ I Introduction
With the advent of Bitcoin and Ethereum [1, 2], we have witnessed the emergence of an increasing number of programmable blockchains  [3, 4]. On these programmable blockchains, developers can write and deploy smart contracts to build various complex decentralized applications (dApps, such as DeFi, NFT, etc. [5, 6]). Among these programmable blockchains, those that share compatible smart contract execution environment with Ethereum (i.e., Ethereum Virtual Machine, a.k.a. EVM [7]) dominate the landscape, accounting for over 90% of the total value locked [8]. With the increasing number of EVM-compatible blockchains and the diversity of dApps running on each chain, the demand for cross-chain dApps has grown significantly [9]. Cross-chain dApps refer to dApps that require cross-chain smart contract invocations (CCSCI) and coordinated execution across multiple blockchains [10].
However, ensuring overall atomicity for the entire cross-chain dApp while efficiently handling CCSCI remains a critical challenge. Consider a classic train-and-hotel problem [11], as illustrated in Figure 1 left. A user wants to book an outbound train ticket (on Train Chain) through a travel agency (on Agency Chain), then book a hotel (on Hotel Chain), followed by a return train ticket (again on Train Chain). The user wants to ensure that the entire series of CCSCI either all succeed or all fail, ensuring overall atomicity. More importantly, maintaining efficiency during CCSCI is crucial. Key considerations include minimizing latency, reducing gas consumption (i.e., monetary cost), and improving concurrency during the CCSCI process by avoiding blocking mechanisms such as global locking. More cross-chain dApp scenarios are discussed in Section VIII.
@@
However, the design of IntegrateX faces several challenges.
Challenge 1: How to efficiently and securely deploy smart contracts across chains. Blindly copying full contracts (logic + state) across chains is gas-intensive and vulnerable to byte-code tampering. We address this by proposing an on-chain/off-chain Hybrid Cross-Chain Smart Contract Deployment Protocol. Specifically, we decouple logic from state, clone only the logic contracts to one target chain (a.k.a. execution chain), and keep state on their original chains. An off-chain byte-code transfer cuts gas costs, while an on-chain hash comparison between the source and target contracts cryptographically proves correctness.
Challenge 2: How to achieve high-concurrency, low-overhead during CCSCI while preserving overall atomicity. During a cross-chain dApp call, multiple states on different chains must be locked, updated, and released as one atomic unit—without significantly throttling performance. We meet this requirement with a 2PC based Cross-Chain Smart Contract Integrated Execution Protocol. It (i) locks only the relevant fine-grained state slices, (ii) aggregates multiple cross-chain state transfer into a single round trip, and (iii) executes the entire dApp call in one transaction before unlocking. This design sustains concurrency and keeps latency and gas low even for deep call graphs.
@@
The remainder of this paper is organized as follows. Section II reviews related work on CCSCI solutions. Section III presents the system model and architecture. Section IV elaborates on the Hybrid Cross-Chain Smart Contract Deployment Protocol. Section V describes the details of the Cross-Chain Smart Contract Integrated Execution Protocol. Section VI analyzes the security of the proposed system. Section VII evaluates the performance of IntegrateX. Section VIII discusses the limitations and future directions. Finally, Section IX concludes the paper.
@@ II Background and Related Work
There are few other solutions that guarantee overall atomicity for the cross-chain dApp with complex CCSCI, as shown in TABLE I. However, they generally suffer from inefficiency  [21, 22, 23]. A complex cross-chain dApp typically involves multiple blockchains and several interdependent application logic components (implemented through smart contracts). In these works, handling such complexity requires multiple rounds of cross-chain execution and cross-chain interaction in sequential order across several involved chains. Specifically, this involves locking relevant states, executing the logic fragment, reaching consensus, transferring intermediate states to the next blockchain responsible for executing the subsequent logic fragment, and finally unlocking and updating the states. As a result, they typically suffer from high latency and poor concurrency handling CCSCI. For example, Robinson et al. introduced GPACT [21] to achieve atomic cross-chain transactions for Ethereum-based blockchains. However GPACT need to lock entire contracts during updates, reducing availability and performance. Additionally, during the execution of cross-chain applications, GPACT requires sequentially waiting for each smart contract call to be executed on its respective blockchain and for the results to be transferred across chains, resulting in significant overhead and high latency. Additionally, Atomic IBC [22] relies on the Cosmos Hub to provide security guarantees for its protocol. This approach is not easily adaptable to other ecosystems (e.g., EVM-compatible blockchains) and may introduce security bottlenecks.
@@
TABLE I: Comparison of Related Works about CCSCI Execution
Strong Atomicity
Compatible with
Existing Blockchain
EVM-compatible Single-round Execution
AtomCI[23],Ivyredaction[39],
GPACT[21],Unity[37]
Atomic IBC[22]
Heterogeneous Paxos[40]
HyperService[36]
SmartSysc[38] read-only
IntegrateX
@@ III System Overview
III-A System Model and Architecture
Infrastructure Components. IntegrateX operates over a network of existing blockchains interconnected via off-chain relayers and on-chain bridging contracts. Figure 2 provides a high-level architecture overview of IntegrateX.
In IntegrateX, there are $n$ (variable) existing blockchains that share the same smart contract execution environment (e.g., EVM-compatible in this paper). These blockchains are developed by their respective projects (e.g., Ethereum, BSC, etc. [2, 44]) and operated by their own blockchain nodes, which handle transaction processing and reach consensus within the blockchain.
IntegrateX also features $m$ (variable) relayers. Relayers are off-chain agents (implemented as independent processes) responsible for listening to events from bridging contracts and relaying information across chains, similar to many mainstream interoperability protocols [22, 15, 45, 46]. Each relayer holds an account (public–private key pair) on each blockchain, and the relayers do not require special privileges; they interact through standard transactions and digitally sign all relayed messages. This requires maintaining a sufficient gas balance on each chain, which can be monitored and refilled as needed via external scripts or relayer management services. To incentivize honest and timely relaying, our system adopts a fee mechanism: users who initiate cross-chain operations are required to pay a cross-chain fee, which is forwarded to the relayer upon successful message delivery. This compensates the relayer for gas costs and operational overhead, while providing an economic incentive aligned with protocol correctness. This model is inspired by, and aligned with, practices in IBC [15], where relayers are independent yet economically motivated.
Additionally, IntegrateX’s bridging smart contracts are deployed on each blockchain (serve as on-chain light clients, like in [15, 22]). The bridging contracts mainly serve to verify cross-chain transaction proofs, register all smart contracts in IntegrateX that are eligible for cross-chain invocation, and emit events externally to initiate cross-chain actions. This transport layer of IntegrateX (relayers + bridging contracts) is similar to that of existing interoperability protocols (e.g., IBC) [15, 22], and provides basic security (ensuring that only valid, agreed-upon cross-chain messages are accepted). However, IntegrateX achieves efficiency and overall atomicity at the application layer, which these other protocols do not.
Application Layer Components. In addition to infrastructure components, we distinguish several roles in the IntegrateX ecosystem’s application layer. Intra-chain dApp providers are developers who create traditional single-chain dApps and deploy the corresponding smart contracts on a blockchain of their choice. These intra-chain providers are not active participants in cross-chain protocols; rather, they supply reusable contract logic that could later be invoked across chains. Cross-Chain dApp providers are developers who compose multiple intra-chain dApps (potentially on different chains) into a cross-chain dApp. A cross-chain dApp typically consists of multiple interrelated smart contracts on different blockchains. The cross-chain provider decides which blockchain will serve as the execution chain (i.e., the single target chain that will host the integrated execution for their dApp’s cross-chain logic) and which other blockchains will be invoked chains supplying state to the execution chain. (In IntegrateX, any EVM-compatible blockchain can be chosen as the execution chain; no custom blockchain is required, as long as the IntegrateX bridging contracts are deployed on it.) Finally, end-users interact with the deployed cross-chain dApps by sending transactions to one of the blockchains (often the execution chain) to trigger the cross-chain functionality.
III-B Threat Model
@@
In IntegrateX, we assume the standard threat model common in cross-chain systems [22, 15, 45, 46, 23, 21]. For each blockchain, the proportion of Byzantine nodes is assumed to be less than the fault tolerance threshold $t$ of the respective blockchain network (e.g., for blockchains using BFT-type consensus protocols under partial-synchronous network, $t=1/3$ [47]). This ensures the safety and liveness of consensus within each blockchain. For the relayers, we make only minimal trust assumptions, assuming that at least one relayer is honest and functioning correctly. This assumption is consistent with those made in many existing secure interoperability protocols [22, 15, 45, 46]. As for the dApp providers and users—who represent the application layer components—we make no specific threat assumptions, similar to most existing works [15, 23, 21]. However, we do discuss common countermeasures for dealing with malicious behavior from these components in Section VIII.
III-C Objective
@@
III-D Overview of the Core Protocols
@@
The operation of IntegrateX consists of two main protocols: Hybrid Cross-Chain Smart Contract Deployment Protocol, and Cross-Chain Smart Contract Integrated Execution Protocol, as illustrated in Figure 2.
Figure 2: The overview of IntegrateX’s core protocols.
Hybrid Cross-Chain Smart Contract Deployment Protocol. To address the high overhead and complexity associated with cross-chain deployment of smart contracts, we design a hybrid deployment protocol that improves efficiency while ensuring security and flexibility. The proposed protocol consists of three core phases: Logic-State Decoupling, Off-Chain Clone and Deployment, and On-Chain Verification, as shown in Figure 2.
The first phase, Logic-State Decoupling, tackles the challenge of redundant data migration in cross-chain scenarios. We devise a set of guidelines that enables dApp developers to separate smart contracts into logic execution contracts and state storage contracts. This approach enables our protocol to clone and deploy only the logic contracts onto the execution chain, while the state-heavy contracts remain in their original locations. In this phase, dApp providers can flexibly develop logic contracts and state contracts according to our defined logic-state decoupling guidelines (Section IV-A).
To further minimize the gas cost and latency involved in cross-chain logic contract deployment, we design an Off-Chain Clone and Deployment mechanism. Instead of performing all operations on-chain, which is resource-intensive, our protocol enables the off-chain cloning and deployment of logic contracts onto the execution chain. This phase is triggered only when dApp providers need to deploy logic contracts from one invoked chain onto the execution chain for execution purposes. The execution chain can be chosen flexibly, and a designated transaction can initiate the deployment process. IntegrateX implements this off-chain deployment technique (Section IV-B) to optimize cost-efficiency.
However, off-chain operations pose integrity risks due to their susceptibility to tampering. To ensure correctness and trustworthiness, we introduce an On-Chain Verification phase. In this phase, bridging smart contracts, which are deployed on both the invoked chain and the execution chain, are used to perform on-chain cross-verification of the logic contract bytecodes between the two chains. This verification mechanism (Section IV-C) guarantees that the deployed contracts are authentic and consistent with their original versions, thereby securing the deployment process against malicious interference.
Through this three-phase protocol design, we effectively reduce deployment overhead while maintaining the integrity and scalability of IntegrateX.
Cross-Chain Smart Contract Integrated Execution Protocol. To achieve high-concurrency, low-overhead and guarantee atomicity during CCSCI, we design an integrated execution protocol based on the two-phase commit (2PC) paradigm, following mainstream practices (Section V-A). The protocol coordinates execution by first locking the relevant contract states across participating invoked chains, transferring these states to a designated execution chain, executing the combined logic, and finally returning the updated states to invoked chains for unlocking and commitment. This ensures that cross-chain transactions either complete entirely or are rolled back in their entirety, preserving consistency.
However, frequent cross-chain state transfers can introduce substantial communication overhead. To address this challenge, we introduce a transaction aggregation mechanism that consolidates multiple cross-chain interactions into a single transaction when applicable, thereby reducing cross-chain communication costs (Section V-B).
Another key issue is the limited concurrency introduced by state locking. In conventional designs, locking entire contract states can hinder parallel transaction processing. To improve concurrency, our protocol incorporates a fine-grained locking scheme (Section V-C). Specifically, we define a set of development guidelines that enable dApp developers to decompose contract states and apply locks at a finer granularity. By isolating and locking only the necessary sub-states, the protocol enables concurrent execution of unrelated operations, significantly enhancing scalability and throughput in multi-chain environments.
Overall, the protocol ensures atomicity, reduces communication overhead, and improves concurrency, thereby enabling reliable and efficient execution of cross-chain dApps.
@@ IV Hybrid Cross-Chain Smart Contract Deployment Protocol
Figure 3: Hybrid Cross-Chain Smart Contract Deployment Protocol.
@@ IV-A1 Logic-State Decoupling Example
function updateState(bytes[] memory args) public returns();
@@
uint256 lock_size;
struct lock_bag;
mapping (uint256 => lock_bag) lockpool;
@@
function updateState(bytes[] memory args) public returns();
@@ IV-B Off-Chain Clone and Deployment
Transmitting contract bytecode on-chain (via contract event) requires a significant amount of gas. To improve the efficiency of cross-chain logic deployment and reduce gas consumption, we propose transferring the contract bytecode via an off-chain solution. The off-chain clone and deployment consists of two phases: the preparation phase and the clone phase. The main process is shown in Figure  3.
Preparation. In this phase, the cross-chain dApp provider needs to obtain the call tree of smart contracts and determine which logic contracts need to be cloned. Similar to existing works [48, 21, 23], this is achieved by first applying static analysis tools such as Slither [49] to extract the call graph of the application. External contract calls can be identified by parsing the call graph, where calls from one contract to another are marked. Static analysis tools can highlight these relationships, helping developers understand cross-contract dependencies. Therefore, from this call graph, the provider constructs the corresponding call tree, which captures the expected invocation paths during a single cross-chain transaction. This process enables the identification of all logic contracts (a superset) that must be cloned to the invoked chain in order to ensure correct and atomic execution. After this, the developer can choose a blockchain to deploy the cross-chain dApp. It is important to note that we offer developers a high degree of flexibility: They can select any blockchain according to their preference. For efficiency, the provider may pick a chain that already hosts some required logic contracts so that those do not need to be cloned again. To reduce costs, they may also choose a blockchain that already hosts some required logic contracts, so that no additional relayer fees are incurred for cloning those contracts again. (Since cross-chain fees are ultimately paid by the provider and compensated to the relayers, this ultimately saves the provider’s expenses.) Once the selection is made, the developer sends a transaction to invoke the bridging contract on the chosen chain. The bridging contract then triggers an event to notify the relayers to initiate the cross-chain deployment. The event includes the ID of the invoked chain and the addresses of the logic contracts $Addr_{\text{L}}$ to be cloned on the invoked chain. This concludes the preparation phase.
Remarks on Call Tree Analysis. Modern static analysis tools provide reliable extraction of contract call graphs and enable accurate construction of call trees. Even in the presence of misconfiguration or adversarial behavior by a dApp provider, the implications of an incorrect call tree are limited. Since the dApp provider is responsible for covering the gas costs of contract cloning, there is a built-in economic disincentive to over- or under-provision the call tree. An incorrect call tree may result in the cloning of unnecessary contracts on the execution chain, thereby increasing storage usage, but it does not introduce any security risk or compromise the atomicity of the cross-chain protocol. Or, a missing logic clone would lead to a runtime exception on the execution chain, effectively aborting the cross-chain transaction. However, our protocol remains atomically safe (no partial commits occur on other chains, but the dApp’s operation would fail). The call tree is not stored on the execution chain. Instead, it is constructed off-chain by the dApp provider during the preparation phase and is used solely as input to the bridging contract when initiating the cross-chain deployment. Only the necessary information—such as the addresses of the logic contracts to be cloned and the identifier of the invoked chain—is packaged and passed as parameters to the bridging contract. This design minimizes on-chain storage and computation overhead, while still providing relayers with the necessary context to coordinate cross-chain deployment.
@@
Remarks on Clone Reliability. For the cross-chain reliability of the protocol, multiple relayers perform the clone process after detecting the bridging contract’s event. Therefore, even if some relayers do not respond to the event, other relayers will ultimately complete the contract clone and deployment. Once one relayer completes the deployment, the bridging contract will trigger an event to stop the others, ensuring that the logic contract will not be deployed multiple times. Furthermore, when a contract is cloned on the execution chain, the deployment transaction will invoke the register function of the bridging contract deployed on the same chain. The bridging contract maintains a mapping from a service identifier (string) to the corresponding logic contract address. Only the first registration for a given identifier will succeed. The deployment transaction will revert if the register fails. This means the contract will not be created on the blockchain, and no bytecode will be stored. This mechanism prevents multiple cloning attempts, as any subsequent invocation of register will fail, causing the entire cloning transaction to fail. As a result, only the first cloning attempt can successfully deploy the contract on the execution chain.
Moreover, if a malicious or overly eager relayer attempts to clone a logic contract without the presence of an event trigger, it must bear the gas cost independently. Relayers are thus economically disincentivized from performing premature cloning, as no reimbursement is provided in the absence of an authorized event. Even if such premature cloning occurs, it does not compromise system correctness; the resulting contracts are merely redundant and remain unused, leading at most to minor storage overhead.